OpenStudy (anonymous):

if the point 3/5,4/5 corresponds to an angle θ in a unit circle what is tan θ

OpenStudy (anonymous):

cos Θ = x coordinate sin Θ = y coordinate tan Θ = y/x

OpenStudy (anonymous):

so I do them two then divide them with tan

OpenStudy (anonymous):

just divide \[\tan \theta=\frac{ 4 }{ 5 }\div \frac{ 3 }{ 5 }\]

OpenStudy (anonymous):

I got a decimal I need another fraction for my answer

OpenStudy (anonymous):

If you used a calculator, it should be able to convert the decimal to a fraction. You could have just multiplied the fractions. \[\tan \theta=\frac{ 4 }{ 5 }\div \frac{ 3 }{ 5 }=\frac{ 4 }{ 5 }\times \frac{ 5 }{ 3 }=\frac{ 4 }{ 3 }\]

OpenStudy (anonymous):

thanks for the help im not allowed a calculator :/ but this really helped thanks peach